OpenStudy (anonymous):

A brick is dropped from a 5 m tall building. At 0.50 s how fast is the brick moving?

OpenStudy (roadjester):

The brick is dropped so it's in free-fall meaning \(\large v_i=0\space\space m/s\). If you apply \(\large v_f=v_i+at\), you'll get your answer.

OpenStudy (anonymous):

would it be 5

OpenStudy (roadjester):

Nope

OpenStudy (anonymous):

k i know the initial v is 0 and vf= 0+1/2 X 9.8?

OpenStudy (roadjester):

? Where did 1/2 come from?

OpenStudy (anonymous):

i have a formula that has 1/2 xgravity

OpenStudy (roadjester):

OH Sorry. I thought t=0.0 s

OpenStudy (roadjester):

My bad

OpenStudy (anonymous):

no prob :)

OpenStudy (roadjester):

Yes if you round 4.9 then it's 5m. Just a word of advice. Conventionally, acceleration is written first, not time.

OpenStudy (anonymous):

0.50+1/2(9.8)?

OpenStudy (roadjester):

No, initial velocity is still 0. \(v_f=at=gt\)

OpenStudy (anonymous):

i have iv =0 vf=0.50

OpenStudy (anonymous):

its 10m/s

OpenStudy (roadjester):

??? Umm Initial velocity=0m/s Acceleration=g=9.8m/s^2 t=0.5s \(v_f = 0m/s+(9.8m/s^2)(0.5s)=4.9m/s\)

OpenStudy (anonymous):

ohhhh thanks i was looking at the wrong formula smt.so basically for this we are trying to the the final velosity

OpenStudy (anonymous):

if i have a question like this can i apply the same formula.... how faw will a brick starting at rest fall freely in 5.0 s?

OpenStudy (roadjester):

If you want to know "how far", then you're looking for position, not velocity which is totally different

OpenStudy (anonymous):

distance deals with how far an object has moved so i can use that instead?

OpenStudy (roadjester):

that's the same thing i just said....

OpenStudy (anonymous):

k so distance is speed x time so 5.0x 0

OpenStudy (roadjester):

\(\huge x_f=x_i+v_it+{1\over 2}at^2\)

OpenStudy (anonymous):

its 49 m